OpenStudy (anonymous):

I need help converting r = 1 / ((5COS(THETA) - 8SIN(THETA)) INTO an equivelent rectangular equation.

OpenStudy (anonymous):

\[x =r \cos \theta\]\[y =r \sin \theta\] this should help solve for cos and sin and put in your equation

OpenStudy (dumbcow):

make the substitutions: \[\frac{x}{8} = 5 \cos \theta\] \[\frac{y}{5} = 8 \sin \theta\] this makes r = 40

OpenStudy (dumbcow):

did you figure it out? \[40 = \frac{1}{\frac{x}{8} - \frac{y}{5}} = \frac{40}{5x-8y}\] \[5x -8y = 1\] \[y = \frac{5x-1}{8}\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

n that case a simpler method of solving it would be just transforming the quotion r(5cos(theta) - 8sin(theta) = 1 : distrubute the r 5rcos(theta) - 8rsin(theta) = 1 : x=rcos(theta), y = rsin(theta) 5x - 8y = 1 solving for y yields the the same result as yours y = (5x - 1) / 8 but I don't think thats the answer

OpenStudy (anonymous):

is 5x - 8y = 1 equal to (x/5) + (y/8) = 1 ?

OpenStudy (dumbcow):

no

OpenStudy (dumbcow):

the polar equation \[r = \frac{1}{5\cos \theta -8 \sin \theta}\] is equal to line \[y = \frac{5x-1}{8}\]

OpenStudy (anonymous):

okay, thank you!

OpenStudy (dumbcow):

yw